OpenStudy (anonymous):

|x − 0.525| = 0.01 solve for x

OpenStudy (nincompoop):

what kind of problem are you having?

OpenStudy (anonymous):

how would i solve the equation to find x

OpenStudy (nincompoop):

what does | x | mean

OpenStudy (anonymous):

the absolute value of x

OpenStudy (nincompoop):

meaning?

OpenStudy (anonymous):

x has to be positive?

OpenStudy (nincompoop):

you are partly correct we were taught that number of value within | | turns into positive

OpenStudy (anonymous):

so x and 0.525 are both positive

OpenStudy (nincompoop):

but if we are dealing with variable such as x, that x can be one of the following: greater than 0 lesser than 0 or equal to zero when it is equal to zero, then we don't really solve anything but if it is less than zero then our equation turns from | x | = -1 (x) if it is greater than zero then it is just positive value or retains its value | x | = x so, to evaluate the expression that you have we have two possible answers -1 (x - 0.525) = 0.01 and the other one is x- 0525 = 0.01 can you solve them now ?

OpenStudy (anonymous):

yes, this was the question i was asked...

OpenStudy (anonymous):

A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the eqation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

would that change it at all?

OpenStudy (nincompoop):

nope

OpenStudy (anonymous):

ok, thank you nin

OpenStudy (nincompoop):

try to solve it so I can check if you understood it clearly

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

if you don't mind, I just don't like to do an half-arsed help. I want to leave the person knowing what I taught.

OpenStudy (anonymous):

yeah, that makes sense

OpenStudy (anonymous):

x = .535?

OpenStudy (nincompoop):

so like I said |x - 0.525 | = 0.01 turns into two equations -1 (x - 0.525) = 0.01 and x - 0.525 = 0.01 first you need to distribute -1 to x then -1 to 0.525 giving you -x + 0.525 = 0.01 solve for x now, then we will worry about the other one after

OpenStudy (nincompoop):

do you know how to evaluate -x + 0.525 = 0.01 ?